OpenStudy (sunnyday):

Solve the equation. Check for extraneous solutions. 9|9 – 8x| = 2x + 3

OpenStudy (sunnyday):

I can get to x=39/37 and x=84/70. From there??

myininaya (myininaya):

I got one of those. I did get x=84/70. But all you need to is to check if both work now. (Also 84/70 can be reduced.)

myininaya (myininaya):

Yep I got both of those. I made a mistake in adding earlier. My bad. Just plug them back in and see you get the same thing on both sides.

OpenStudy (sunnyday):

How do I do that using fractions?

myininaya (myininaya):

Just like you would with integers.

OpenStudy (sunnyday):

9(9-8 (39/37))=2(39/37)+3 ??

myininaya (myininaya):

\[|9-8 \cdot \frac{39}{37}|=\frac{2}{9} \cdot \frac{39}{37}+\frac{3}{9} \text{ Does this check out?}\]

OpenStudy (sunnyday):

OK. yes

myininaya (myininaya):

Then you would do the same for x=84/70 (=6/5 if not mistaken) \[|9-8 \cdot \frac{6}{5}|=\frac{2}{9} \cdot \frac{6}{5}+\frac{3}{9} \text{ Does this check out?}\]

myininaya (myininaya):

Oh and your equation is fine to use. I just divided both sides by 9 before checking.

OpenStudy (sunnyday):

Yes, I caught that.

OpenStudy (sunnyday):

Do you recall the order of operations as far as mly, d, parenthesis, exponents ....??

OpenStudy (sunnyday):

wow - add, multiply is what that should say!

myininaya (myininaya):

Do the operations inside the grouping symbols... Like here we have the absolute value function which includes operations like multiplication and subtraction. So in the absolute value part you would do the multiplication first and then the subtraction. |9-8(39/37)|=|9-312/37| Remember to subtract (or add) fractions we need to find a common denominator.

myininaya (myininaya):

You simplify both sides using order of operations. Then you see if both sides match.

OpenStudy (sunnyday):

312 came from where?

myininaya (myininaya):

Multiplying 8 and 39/37. I'm following the order of operations.

OpenStudy (sunnyday):

ok. I was onside o the equation trying to figure out where 312 came from.

OpenStudy (sunnyday):

Thank you.

myininaya (myininaya):

You got it from there?

OpenStudy (sunnyday):

not sure.

OpenStudy (sunnyday):

seems like a lot of work. plugging the fractions in. I'm not good with fractions. trying.

myininaya (myininaya):

multiplying them or adding them?

OpenStudy (sunnyday):

both :/

myininaya (myininaya):

\[\frac{a}{b} \cdot \frac{c}{d}=\frac{a \cdot c}{b \cdot d} ; \frac{a}{b} \pm \frac{c}{d}=\frac{a \cdot d \pm c \cdot b}{b \cdot d}\]

OpenStudy (sunnyday):

so would you first multiply 2/9 and 39/37?

OpenStudy (sunnyday):

having the denominator 333?

OpenStudy (sunnyday):

78/333 +3/9=

myininaya (myininaya):

For the other side of the equation? Yes we have \[\frac{2}{9} \cdot \frac{39}{37} +\frac{1}{3}\] mult first \[\frac{36}{111}+\frac{1}{3}\] Yes the denominator should be 333 \[\frac{36(3)+111(1)}{3(111)}\]

OpenStudy (sunnyday):

how did u get 111? 9(37) is 333

OpenStudy (sunnyday):

LCF?

myininaya (myininaya):

I reduced the fraction during multiplication.

OpenStudy (sunnyday):

I got 78/333 + 1/3

OpenStudy (sunnyday):

2(39)/9(37)

OpenStudy (sunnyday):

+1/3 of course

myininaya (myininaya):

\[\frac{2}{9}\cdot \frac{39}{37}=\frac{2}{3 \cdot 3} \cdot \frac{3 \cdot 13}{37}=\frac{2}{3 \cdot \cancel{3}} \cdot \frac{\cancel{3} \cdot 13}{37}=\frac{2 \cdot 13}{3 \cdot 37}\]

myininaya (myininaya):

26/111 (and I made a type-o above)

myininaya (myininaya):

This keyboard has sticky buttons :(

OpenStudy (sunnyday):

OK. 78/333 (my answer) if reduced is 26/111. We got the same answer. Now does that equal /9-312/37/??

myininaya (myininaya):

Did you add the 1/3 yet?

OpenStudy (sunnyday):

oops. no.

OpenStudy (sunnyday):

so 26/111+1/3

myininaya (myininaya):

\[\frac{2}{9} \cdot \frac{39}{37}+\frac{1}{3}=\frac{26}{111}+\frac{1}{3}=\frac{26(3)+1(111)}{111(3)}\]

OpenStudy (sunnyday):

189/333

OpenStudy (sunnyday):

so, now is = to/9-312/37/?

myininaya (myininaya):

So how do you subtract 312/37 from 9?

OpenStudy (sunnyday):

I would make it a mixed number but I've been to use fractions. so, put a 1 under the 9??

myininaya (myininaya):

9(37/37)-312/37 So we can have the same denominator